OpenStudy (anonymous):

If f(x) = 3x2 + 3x and g(x) = x2 - 5x + 2. What is f(x) + g(x)? Answer 3x4 - 15x2 + 2 3x2 - 15x + 2 4x2 - 8x + 2 4x2 - 2x + 2

OpenStudy (campbell_st):

just add them 3x^2 + 3x + x^2 -5x + 2 = 4x^2 -2x + 2

OpenStudy (anonymous):

i dont understand how to do this problem

Directrix (directrix):

f(x) + g(x) means the sum of the terms of f(x) and g(x) 3x^2 + 3x + x^2 - 5x + 2 = 4 x^2 - 2x + 2 f(x) + g(x) = 4 x^2 - 2x + 2

OpenStudy (campbell_st):

the question of asking to just add the 2 functions....f(x) and g(x) f(x) + g(x)